OpenStudy (anonymous):

(7x – 6y) + (3x + 3y – 3)

OpenStudy (anonymous):

=7x+−6y+3x+3 Combine Like Terms: =7x+−6y+3x+3 =(7x+3x)+(−6y)+(3) =10x+−6y+3 Answer: =10x−6y+3

OpenStudy (zephyr141):

@REBAATONDO you forgot the positive 3y and the signs are switched for the 3 so it should be 10x-3y-3

OpenStudy (anonymous):

oh im sorry @zephyr141